(8) Solve the following equations. Remember to check your solutions. (a) V6x + 7 = x+2 (b) VX-3 + 5 = x
Answer #1

Cou [6x+7= 2+2 squating on both sides (562+7) (x+2) 62+7 +91-9=6 x 32+2-3 =0 - *(x-3) +(2-3) - = (2+1) (2-3) =0 - 25-1 ol 2 =